You are driving home from school steadily at 93 km/h for 161 km. It then begins to rain and you slow to 54 km/h. You arrive home after 6 h. How far is your hometown from school?
1
Expert's answer
2020-09-01T10:37:23-0400

The 6 hours consist of 161 km passed at 93 km/h and some distance passed at 54 km/h:


"6=\frac{161}{93}+\frac{x}{54},\\\space\\
x=230.5\text{ km}."

The total distance is


"D=161+230.5=391.5\text{ km}."